# Tight Bounds for Sparsifying Random CSPs

### v1

@@ 1.3.1 A Crude Sparsifier
To see the lower bound, we consider the sets $D_{1},\dots D_{r}$ exactly realizing the $\mathbf{AND}$ of arity $c$, and we let $a_{1}a_{2}\dots a_{r}$ be the unique satisfying assignment in $R\cap(D_{1}\times D_{2}\times\dots\times D_{r})$. We let $T\subseteq[r]$ denote the indices where the $D_{i}$’s are of size $2$, i.e., $T=\{i\in[r]:|D_{i}|=2\}$. With this, we can describe our set of assignments $X$ that we will use to derive the lower bound. For each assignment, for every $i\in T$, we choose a special distinguished vertex $x^{(i)}_{j^{*}_{i}}\in V_{i}$, and set $x^{(i)}_{j^{*}_{i}}=a_{i}$, while for every other variable in $V_{i}$, we set it to $\neg a_{i}$. For the remaining indices $i\notin T$, we set all the variables to $a_{i}$. In this way, an assignment is exactly described by its distinguished vertices in $V_{i}:i\in T$.
Now intuitively, for each assignment $\psi\in X$, the only satisfied constraints will be exactly those that contain all of the distinguished vertices for that assignment. This effectively creates a partition of the constraints into $n^{c}$ disjoint classes, where for the class $X$ of assignments, no constraint is simultaneously satisfied by two different assignments. Thus, any sparsifier is forced to keep one satisfied constraint for each assignment, leading to an $\Omega(n^{c})$ lower bound.
@@
constraints to survive. By a Chernoff bound then, a random subsampling will preserve the fraction of satisfied assignments for assignment $\psi$ with probability $1-2^{-\Omega(n)}$, which is sufficiently large to survive a union bound over all $2^{O_{r}(n)}$ possible assignments $\psi$. Thus, our subsampling at rate roughly $\frac{1}{\varepsilon^{2}\cdot n^{r-c-1}}$ will constitute a sparsifier for $C$ with high probability.
@@
However, in order to shave of the extra +1+1 in the exponent of the sparsifier size, we need to take a more careful look at the structure of the satisfied constraints.
@@ 1.3.2 Attempts at Improving the Upper Bound
To show that such a sampling scheme would work in our context, a counting bound analogous to [Kar93]’s would require that the number of assignments $\psi\in\{0,1\}^{nr}$ for which the number of satisfying satisfied constraints is positive and at most $\lambda\cdot n^{r-c}$ is at most $n^{O(\lambda)}$. Unfortunately, it turns out that such a counting bound is not true. Indeed, we can consider the simple relation $R=\{000,001\}\subseteq\{0,1\}^{3}$, where the largest arity $\mathbf{AND}$ is of size $c=2$. Now, consider the family of assignments $X=\{\psi\in\{0,1\}^{3n}:\psi_{1:2n}=0\circ 1^{n-1}\circ 0\circ 1^{n-1}\}$. For each such assignment, exactly $n$ constraints will be satisfied, corresponding to the constraints which contain variables $x_{1}$ and $x_{n+1}$, and any choice of variable in $V_{3}$. However, because we allow the assignments in $X$ to vary arbitrarily in $V_{3}$, $|X|=2^{n}$, and so we have an exponential number of assignments satisfying $n$ constraints, thereby providing a strong refutation to our smooth counting bound.
@@ 1.3.3 Decomposing Relations Through Irrelevant Coordinates
Going forward, we restrict our attention to assignments $\psi\in\{0,1\}^{nr}$ which are $0^{r}$-dominant, meaning that for every $i\in[r]$, $|\psi^{-1}(0)\cap V_{i}|\geq n/2$. We will assume that $0^{r}$ is one such extreme tuple for the relation $R$. Note that if $0^{r}$ were not extreme, then this means that $0^{r}$ is distance $\leq c-1$ from the relation $R$. So, for any assignment $\psi$ such that for every $i\in[r]$, $|\psi^{-1}(0)\cap V_{i}|\geq n/2$ (i.e., an assignment where in each part, the most common value is a 0), such an assignment would actually satisfy at least $(n/2)^{n-r+1}$ constraints, a factor $n$ improvement over our crude analysis in 1.6. Thus, the more interesting case is when a tuple (like $0^{r}$) is distance exactly $c$ from $R$.
Now, with this established, we consider the tuple $t=0^{r}$, and assume it is extreme for $R$. As mentioned above, let us assume that $0^{r}$ has distance $c$ from $R$. We consider all possible sets $S\in\binom{[r]}{c}$ such that $0^{r}\oplus e_{S}=e_{S}\in R$ (corresponding to which bits must be flipped). For each index $i\in[r]$, we look at whether there exists an $S$ such that $e_{S}\in R$ and $i\in S$. If no such set exists, we say that the coordinate $i$ is irrelevant, as intuitively, this coordinate never acts as a witness to the tuple $0^{r}$ being distance $c$ from $R$.
In fact, there is an even stronger sense in which so-called irrelevant coordinates are indeed irrelevant:
@@
As further motivation for the above proposition, we introduce the notion of a decomposable relation. Again, we consider the case when $t=0^{r}$ (assuming it is extreme), and we let $I$ denote the set of irrelevant coordinates for $t$. We say that $R$ is decomposable if $R=Q\times\{0,1\}^{I}$, where $I$ is the set of irrelevant coordinates and $Q\subseteq\{0,1\}^{\bar{I}}$. The benefit of this characterization is that any decomposable predicate satisfies a smooth counting bound of our desired form, when we consider the “codeword” perspective of the CSP: i.e., for an assignment $\psi\in\{0,1\}^{nr}$, we consider the corresponding vector in $\{0,1\}^{m}$ (where $m$ is the number of constraints) which is the indicator vector of the satisfied constraints under $\psi$.
@@
Using our assumption that $|\psi^{-1}(0)\cap V_{i}|\geq n/2$, this then evaluates to $\mathrm{sat}_{R}(\psi)\geq(n/2)^{r-c}\cdot|\psi^{-1}(1)\cap V_{i}|$. In particular, if $\psi$ satisfies fewer than $\lambda\cdot n^{r-c}$ constraints, then it must be the case that $s_{i}=O_{r}({\lambda})$, and so there are at most $\binom{n}{O_{r}({\lambda})}=n^{O_{r}(\lambda)}$ choices for how to choose the positions of $1$’s in the $i$th part of the assignment $\psi$.
@@ 1.3.4 Sandwiching to Create Sparsifiers
In the previous section we established a counting bound (and thus sparsifiability) for so-called decomposable relations. However, in general there is no guarantee that a relation $R$ forms this type of product structure with the irrelevant coordinates, as the above proposition only guarantees that extreme tuples can be arbitrarily extended with irrelevant coordinates while maintaining satisfaction. So, it is possible for instance, that there are tuples of weight $c+1$ in $R$ which do depend on the irrelevant coordinates, and therefore violate our ability to “group” assignments by only their relevant coordinates, as now their corresponding codewords will be different. How then can we still guarantee the sparsifiability of such CSPs?
@@
1.3.5 Barriers for Arbitrary Domains and Uniform CSPs
The above analysis is very specifically tailored to the setting of Boolean relations, as our notion of irrelevance depends on exactly which coordinates are $1$ in a tuple of Hamming weight $c$ in $R$. To generalize the argument to relations $R\subseteq D^{r}$, we also need to generalize our notion of irrelevance to pairs of coordinates and symbols. That is to say, instead of considering a set of irrelevant coordinates $I\subseteq[r]$, we consider $I\subseteq[r]\times D$. As before, we can follow a similar path to show a type of “product structure” with these irrelevant pairs, followed by a sandwiching argument to establish sparsifiability.
The generalization to uniform CSPs turns out to be the most technically challenging. In this case, the relation $R$ gets symmetrized (as all permutations of each relation are equally likely) and valued by default, and so instead of caring about the largest $\mathbf{AND}$ in $R$, we are concerned with a type of “Pareto frontier” tuple in $R$. Specifically, we fix a domain element $d$, and look at tuples $t$ in $R$ such that there is no other tuple $t^{\prime}$ satisfying that for every symbol $e\in D-\{d\}$, $_{e}(t^{\prime})\geq\#_{e}(t)$ (where $_{e}(t)$ means the number of occurrences of $e$ in $t$).
At a high level, the goal is unchanged from the $r$-partite setting: i.e., to establish a smooth counting bound and then pair it with sandwiching arguments to conclude the sparsifiability of the CSP. However, the translation from an $r$-partite CSP to a uniform CSP leads to a different characterization of sparsifiability which admits a much trickier edge case: for illustration, consider the two uniform CSPs, one with relation $R=\{0,1\}^{r}$, and another with relation $R=\{0^{r},1^{r}\}$. In both cases, every assignment $\psi\in\{0,1\}^{n}$ satisfies $\Omega(n^{r})$ constraints, and so it would be tempting to argue that the appropriate sampling rate for creating a sparsifier should be roughly $\frac{1}{n^{r}}$. However, it turns out that this sampling rate only works for creating sparsifiers of $R=\{0,1\}^{r}$, and for $R=\{0^{r},1^{r}\}$, any sparsifier needs $\Omega(n)$ constraints. In the Boolean case, this turns out to be the only “tricky” instance to deal with, but when generalizing to arbitrary domains, this type of “uniform” vs “non-uniform” sub-instance becomes much more pathological, leading to the somewhat counter-intuitive classifications that we derive in this work. For brevity of the technical overview, we delay a full discussion of these details to Section 7.
@@ 1.4 Open Questions
    What is the sparsifiability of random CSPs using a relation $R\subseteq D^{r}$ when $r$ grows with the number of variables $n$ (that is to say, when $r$ is no longer constant)?
@@
    What is the spectral sparsifiability of random CSPs (as defined in [KPS25b])?
@@
    Can we understand the sparsifiability of CSPs in semi-random models? As an example, consider an instance in which half of the constraints are chosen by an adverary and half of the constraints are chosen uniformly at random. Which half of the instance will “dominate” the asymptotics of the size of the optimal sparsifier?
@@ 2.1 Notation
Note that throughout this work, we will often refer to a CSP $C$, with the understanding that we are referring to a CSP instance $C$. For a CSP $C$ and an assignment $x$, we will use the notation $\mathrm{sat}_{C}(x)$ to denote the number of satisfied assignments in $C$ under the assignment $x$. Additionally, we will often use the notion of codewords for an assignment $x$:
@@ 3 Sparsifiability of the Complete, Uniform Instance with Boolean Relations
As above, let $c=\max(\operatorname{wt}(a_{\min}),r-\operatorname{wt}(a_{\max}))$, and suppose $c\geq 1$. Now, if $c=\operatorname{wt}(a_{\min})$, then let $b=0$, and otherwise, let $b=1$. For our lower bound, we consider the set of assignments
@@ 4 Sparsifiability of the Complete, $r$-Partite Instance with Boolean Relations
With this, we present beow our characterization of the sparsifiability of complete, $r$-partite, Boolean CSPs. Note that although the notion of non-redundancy can be defined for the complete instance $C$, non-redundancy only governs the worst sparsifiability of any sub-instance of $C$. In general, by including all constraints, we can actually get smaller size sparsifiers:

### v2

@@ 1.3 Technical Overview
1.3.1 Intuition
In order to derive all of the above theorems, our results rely on several new definitions, techniques, and decompositions for CSPs. As mentioned above, the best known characterizations of CSP sparsification like [BG25] (and for other classes of sparsifiers more generally, like [BKLM25]) struggle when dealing with instance-optimal sparsification, as these methods can only guarantee sparsifier sizes equal to the sparsifiability of the worst-case sub-instance. For instance, when invoking the main result of [BG25] on a complete directed graph, the sparsifier size will only be guaranteed to be $O(n^{2})$, whereas we know from the previous discussion that sparsifiers of size $\widetilde{O}(n)$ exist. It turns out that this same phenomenon presents us with difficulties in our current setting as well. Indeed, one of the main barriers to optimal sparsification of random CSPs is that there can be too many assignments who satisfy a small number of constraints, (and thus random sampling + a naive Chernoff and union bound does not work for sparsification) as a result of a small subset of “unluckily” sampled constraints.
However, for a random CSP instance $C$ with relation $R$, we provide a general structural theorem that shows that this excess of assignments satisfying a small number of constraints can be remedied by instead considering a “sandwich CSP”, whereby we replace the relation $R$ with two relations $R_{0}\subseteq R\subseteq R_{1}$ of a specific form. [^9] Note that these types of sandwiches are also studied phenomena in graph theory, under the label of graph sandwich problems [GKS95]. We show that these new CSPs $C_{0}$, $C_{1}$ that result from replacing $R$ with $R_{0}$ and $R_{1}$ now are sparsifiable by a random sample to size equal to our classification theorems above. Thus, for a given assignment, if it is the case that $C_{0}$ and $C_{1}$ have approximately the same number of satisfied constraints, then the simple fact that $R$ is sandwiched (i.e., $R_{0}\subseteq R\subseteq R_{1}$) implies that the original CSP $C$ is also tightly sandwiched in between $C_{0}$ and $C_{1}$, and so a random sample of $C$ will also be a sparsifier of $C$ with high probability. Otherwise, if $C_{0}$ and $C_{1}$ have a vastly differing number of satisfied constraints for a given assignment, we show that (using the structure of our specific sandwiching relations $R_{0},R_{1}$) this is because $C$ itself must have many satisfied constraints for this assignment. In this latter case, we are no longer dealing with the “low-weight” regime, and can now afford a Chernoff bound and union bound argument.
Designing these sandwiching predicates itself relies on a delicate analysis of the influence of specific variables in our relation $R$. In the $r$-partite setting, this relies on notions of relevant and irrelevant coordinates, as we discuss in the deeper technical overview (Section 1.3.4), and in the more difficult arbitrary domain, uniform CSP setting, relies on more involved notions of “plentifulness” of the relation $R$, which we omit from the discussion here for brevity.
Nevertheless, these methods provide a concrete framework for circumventing the unsparsifiability of sub-instances in a CSP by analyzing CSPs with carefully chosen, close relations where these sub-instances are no longer unsparsifiable. We view it as a very promising future direction to use these techniques and this broader paradigm to derive “instance-optimal” characterizations of sparsifiability; i.e., to say for a fixed CSP instance $C$ what the optimal sparsifier size is.
In the following section, we explain these techniques in more detail, by deriving (part of) the characterization in the significantly simpler Boolean setting.
1.3.2 A Crude Sparsifier
@@
To see the lower bound, we consider the sets $D_{1},\dots D_{r}$ exactly realizing the $\mathbf{AND}$ of arity $c$, and we let $a_{1}a_{2}\dots a_{r}$ be the unique satisfying assignment in $R\cap(D_{1}\times D_{2}\times\dots\times D_{r})$. We let $T\subseteq[r]$ denote the indices where the $D_{i}$’s are of size $2$, i.e., $T=\{i\in[r]:|D_{i}|=2\}$. Now, we consider the set of assignments $X\subseteq\{0,1\}^{nr}$ which, for indices $i\in[r]-T$ always take value $a_{i}$ for variables in $V_{i}$, and for indices $i\in T$ take value $a_{i}$ in exactly one variable in $V_{i}$, which we call the distinguished variable, and are otherwise given value $1-a_{i}$.
Now intuitively, for an assignment $\psi\in X$, the only satisfied constraints will be exactly those that contain all of the distinguished variables for that assignment. For each $V_{i}:i\in T$ there are $n$ choices of the distinguished variable, so this effectively creates a partition of the constraints into $n^{c}$ disjoint classes where no assignment $\psi\in X$ simultaneously satisfies constraints from two different classes. Thus, for any sparsifier to ensure that every assignment in $X$ still satisfies at least one constraint, it must be the case that the sparsifier retains at least one constraint from each class, thus yielding an $n^{c}$ lower bound.
@@
constraints to survive. By a Chernoff bound then, a random subsampling will preserve the fraction of satisfying assignments for assignment $\psi$ with probability $1-2^{-\Omega(n)}$, which is sufficiently large to survive a union bound over all $2^{O_{r}(n)}$ possible assignments $\psi$. Thus, our (reweighted) subsampling at rate roughly $\frac{1}{\varepsilon^{2}\cdot n^{r-c-1}}$ will constitute a sparsifier for $C$ with high probability.
@@
However, in order to shave off the extra +1+1 in the exponent of the sparsifier size, we need to take a more careful look at the structure of the satisfied constraints.
@@ 1.3.3 Attempts at Improving the Upper Bound
To show that such a sampling scheme would work in our context, a counting bound analogous to [Kar93]’s would require that the number of assignments $\psi\in\{0,1\}^{nr}$ for which the number of satisfied constraints is positive and at most $\lambda\cdot n^{r-c}$ is at most $n^{O(\lambda)}$. Unfortunately, it turns out that such a counting bound is not true. Indeed, we can consider the simple relation $R=\{000,001\}\subseteq\{0,1\}^{3}$, where the largest arity $\mathbf{AND}$ is of size $c=2$. Now, consider the family of assignments $X=\{\psi\in\{0,1\}^{3n}:\psi_{1:2n}=0\circ 1^{n-1}\circ 0\circ 1^{n-1}\}$. For each such assignment, exactly $n$ constraints will be satisfied, corresponding to the constraints which contain variables $x_{1}$ and $x_{n+1}$, and any choice of variable in $V_{3}$. However, because we allow the assignments in $X$ to vary arbitrarily in $V_{3}$, the size of $|X|$ is $2^{n}$, and so we have an exponential number of assignments satisfying $n$ constraints, thereby providing a strong refutation to our smooth counting bound.
@@ 1.3.4 Decomposing Relations Through Irrelevant Coordinates
Going forward, we restrict our attention to assignments $\psi\in\{0,1\}^{nr}$ which are $0^{r}$-dominant, meaning that for every $i\in[r]$, $|\psi^{-1}(0)\cap V_{i}|\geq n/2$. We will assume that $0^{r}$ is one such extreme tuple for the relation $R$. Note that if $0^{r}$ were not extreme, then this means that $0^{r}$ is distance $\leq c-1$ from the relation $R$. So, for any assignment $\psi$ such that for every $i\in[r]$, $|\psi^{-1}(0)\cap V_{i}|\geq n/2$ (i.e., an assignment where in each part, the most common value is a 0), such an assignment would actually satisfy at least $(n/2)^{r-c+1}$ constraints, a factor $n$ improvement over our crude analysis in 1.6. Thus, the more interesting case is when a tuple (like $0^{r}$) is distance exactly $c$ from $R$.
Now, we look at the entire set of tuples in $R$ which are at distance exactly $c$ from $t=0^{r}$: formally, we consider all possible sets $S\in\binom{[r]}{c}$ such that $0^{r}\oplus e_{S}=e_{S}\in R$ (corresponding to which bits must be flipped). For each index $i\in[r]$, we look at whether there exists an $S$ such that $e_{S}\in R$ and $i\in S$. If no such set exists, we say that the coordinate $i$ is irrelevant, as intuitively, this coordinate never acts as a witness to the tuple $0^{r}$ being distance $c$ from $R$.
In fact, there is an even stronger sense in which so-called irrelevant coordinates are irrelevant:
@@
As further motivation for the above proposition, we introduce the notion of a decomposable relation. Again, we consider the case when $t=0^{r}$ (assuming it is extreme), and we let $I$ denote the set of irrelevant coordinates for $t$. We say that $R$ is decomposable if $R=Q\times\{0,1\}^{I}$, where $I$ is the set of irrelevant coordinates and $Q\subseteq\{0,1\}^{\bar{I}}$. The benefit of this characterization is that any decomposable predicate satisfies a smooth counting bound of our desired form, when we consider the “codeword” perspective of the CSP: i.e., for an assignment $\psi\in\{0,1\}^{nr}$, we associate it with the corresponding vector in $\{0,1\}^{m}$ (where $m$ is the number of constraints) which is the indicator vector of the satisfied constraints under $\psi$. Morally, this perspective treats two assignments $\psi_{1}$ and $\psi_{2}$ as being the same if the sets of constraints satisfied by $\psi_{1}$ and $\psi_{2}$ are the same.
@@
Using our assumption that $|\psi^{-1}(0)\cap V_{i}|\geq n/2$, this then evaluates to $\mathrm{sat}_{R}(\psi)\geq(n/2)^{r-c}\cdot|\psi^{-1}(1)\cap V_{i}|$. In particular, if $\psi$ satisfies fewer than $\lambda\cdot n^{r-c}$ constraints, then it must be the case that $|\psi^{-1}(1)\cap V_{i}|=O_{r}({\lambda})$, and so there are at most $\binom{n}{O_{r}({\lambda})}=n^{O_{r}(\lambda)}$ choices for how to choose the positions of $1$’s in the $i$th part of the assignment $\psi$.
@@ 1.3.5 Sandwiching to Create Sparsifiers
In the previous section we established a counting bound (and thus sparsifiability) for so-called decomposable relations. However, in general there is no guarantee that a relation $R$ forms this type of product structure with the irrelevant coordinates, as the above proposition only guarantees that extreme tuples can be arbitrarily extended with irrelevant coordinates while maintaining satisfaction. So, it is possible for instance, that assignments $\psi_{1}$ and $\psi_{2}$ do agree on constraints satisfied by extreme tuples, but nevertheless satisfy different sets of constraints overall due to the “lower order” contributions from tuples of weight $c+1$ in $R$. How then can we still guarantee the sparsifiability of such CSPs?
@@
While these techniques ultimately suffice for deriving a characterization of the sparsifiability of $r$-partite Boolean CSPs, we dedicate the majority of this work towards the challenges that arise in arbitrary domains, valued CSPs, and the uniform setting. For the sake of brevity, we postpone this discussion to the main body of the work.
@@ 1.4 Open Questions
    Can the sandwiching and decomposability techniques be used to derive a general instance-optimal characterization of sparsifiability? I.e., to characterize for any specific CSP instance $C$ what the optimal sparsifier size is for $C$. Note that, as mentioned before, existing characterizations [BG25] are always lower bounded by the worst-case sparsifiability of any sub-instance of $C$ which is not in general equal to the sparsifiability of $C$ as a whole.
@@
    What is the sparsifiability of random CSPs using a relation $R\subseteq D^{r}$ when $r$ grows with the number of variables $n$ (that is to say, when $r$ is no longer constant)? Note that in our results, our sparsifier sizes depend (often exponentially) on the number of variables $r$ and the size of the domain $|D|$. Thus, even for relations $R\subseteq\{0,1\}^{\mathrm{polylog}(n)}$, our characterizations lose their tightness.
@@
    What is the spectral sparsifiability of random CSPs (as defined in [KPS25b])? Here, instead of considering assignments in $\{0,1\}^{r}$, one instead considers continuous-valued assignments in $0,1]^{r}$. For a single constraint $c$ in a CSP, the value of $c$ on assignment $x\in[0,1]^{r}$ is then defined to be $\Pr_{\theta\sim[0,1]}[c(x^{(\theta)})=1]^{2}$, where $x^{(\theta)}$ refers to the rounded version of $x$, where every entry below $\theta$ gets sent to 0, and every entry larger than $\theta$ gets sent to $1$. As shown in [KPS25b], this notion generalizes spectral sparsification in graphs and hypergraphs, and thus one can naturally ask whether our characterizations hold for this stronger notion of sparsification.
  4. 4.
    Can we understand the sparsifiability of CSPs in semi-random models? As mentioned above, one of our primary contributions is to give a framework for reasoning about how the worst-case sparsifiability of a sub-instance of a given CSP affects the global sparsifiability of the entire CSP. One way to further test this framework is to instead consider a model where $m$ constraints are generated at random, and then for some fraction $p\in(0,1)$, an additional $p\cdot m$ constraints are chosen by an adversary. As a function of $p,R,m,n$, what then governs the sparsifiability of the resulting instance? Intuitively, this question asks how large an adversarial worst-case sub-instance must be before it “dominates” the asymptotics of the optimal sparsifier size.
@@ 2.1 Notation
Note that throughout this work, we will often refer to a CSP $C$, with the understanding that we are referring to a CSP instance $C$. For a CSP $C$ and an assignment $x$, we will use the notation $\mathrm{sat}_{C}(x)$ to denote the number of satisfied constraints in $C$ under the assignment $x$. Additionally, we will often use the notion of codewords for an assignment $x$:
@@ 3 Sparsifiability of the Complete, Uniform Instance with Boolean Relations
As above, let $c=\max(\operatorname{wt}(a_{\min}),r-\operatorname{wt}(a_{\max}))$, and suppose $c\geq 1$. Now, if $c=\operatorname{wt}(a_{\min})$, then let $b=1$, and otherwise, let $b=0$. For our lower bound, we consider the set of assignments
@@ 4 Sparsifiability of the Complete, $r$-Partite Instance with Boolean Relations
With this, we present below our characterization of the sparsifiability of complete, $r$-partite, Boolean CSPs. Note that although the notion of non-redundancy can be defined for the complete instance $C$, non-redundancy only governs the worst sparsifiability of any sub-instance of $C$. In general, by including all constraints, we can actually get smaller size sparsifiers: